OpenStudy (anonymous):

Find the specified vector or scalar. u = -4i + 1j and v = 4i + 1j; Find ||u + v||

OpenStudy (anonymous):

A. sqrt34 B. 8 C. 5 D. 2

ganeshie8 (ganeshie8):

u = -4i + 1j v = 4i + 1j ------------------

ganeshie8 (ganeshie8):

to add them both, add each component

ganeshie8 (ganeshie8):

u = -4i + 1j v = 4i + 1j ------------------ u+v = 0i + 2j

ganeshie8 (ganeshie8):

u knw how to take magnitude of that vector right ?

ganeshie8 (ganeshie8):

use below formula :- \(\large \mathbb{|| xi + yj || = \sqrt{x^2+y^2}}\)

ganeshie8 (ganeshie8):

\(\large \mathbb{|| u + v || = ||0i + 2j|| = \sqrt{0^2+2^2} = \sqrt{0+4} = \sqrt{4} = 2}\)

OpenStudy (anonymous):

That makes sense!! :D Thank you!!

ganeshie8 (ganeshie8):

good :)